OpenStudy (daggiemib):

For a community breakfast there should be at least three times as much regular coffee as decaffeinated coffee. A total of ten gallons is sufficient for the breakfast. Write and graph a system of inequalities to model the problem.

OpenStudy (anonymous):

Here's one inequality (r=regular, d=decaf). See if you can figure out the other. r (greater than or equal to) 3d

OpenStudy (daggiemib):

\[y \ge 3x\] & \[x+y \ge 10\] turns into \[y \ge -x+10\] then i graph them

OpenStudy (anonymous):

correct! nice job